OpenStudy (anonymous):

Find the common ratio of the sequence. 224, 56, 14, 3.5, . . .

OpenStudy (anonymous):

r=0.25

OpenStudy (anonymous):

56/224 = 0.25 14/56=0.25 3.5/14=0.25